### v1

# Lifelong Knowledge Editing for Vision Language Models with Low-Rank Mixture-of-Experts

Abstract
Model editing aims to correct inaccurate knowledge, update outdated information, and incorporate new data into Large Language Models (LLMs) without the need for retraining. This task poses challenges in lifelong scenarios where edits must be continuously applied for real-world applications. While some editors demonstrate strong robustness for lifelong editing in pure LLMs, Vision LLMs (VLLMs), which incorporate an additional vision modality, are not directly adaptable to existing LLM editors. In this paper, we propose LiveEdit, a Lifelong vision language model Edit to bridge the gap between lifelong LLM editing and VLLMs. We begin by training an editing expert generator to independently produce low-rank experts for each editing instance, with the goal of correcting the relevant responses of the VLLM. A hard filtering mechanism is developed to utilize visual semantic knowledge, thereby coarsely eliminating visually irrelevant experts for input queries during the inference stage of the post-edited model. Finally, to integrate visually relevant experts, we introduce a soft routing mechanism based on textual semantic relevance to achieve multi-expert fusion. For evaluation, we establish a benchmark for lifelong VLLM editing. Extensive experiments demonstrate that LiveEdit offers significant advantages in lifelong VLLM editing scenarios. Further experiments validate the rationality and effectiveness of each module design in LiveEdit.  [^1] Source codes are provided in the supplementary material and will be released upon acceptance.
1 Introduction
Large language models (LLMs) have become key techniques for text generation in NLP [1, 2, 3]. Benefiting from vision-language pre-training and pure LLMs, Vision-LLMs (VLLMs) are capable of generating text responses based on images and text [4, 5, 6, 7]. However, outdated or erroneous built-in knowledge can undermine the value of these models. To avoid the costly retraining of large-scale parameters, model editing aims to adapt models by adjusting a small number of parameters to update specific knowledge. This plays a critical role in areas such as privacy protection [8, 9], detoxification [10, 11], bias reduction [12, 13, 14], and hallucination correction [15, 16].
Figure 1: Lifelong VLLM Editing. In this scenario, the edited VLLM is required to correctly respond to queries involving the edited data within the generalization domain, while maintaining consistent responses in locality domains. The top left shows test cases where Rel., T-Gen./M-Gen., T-Loc., and M-Loc. denote reliability, text/modal generality, and text locality, respectively. The bottom right illustrates the responses of an effectively edited VLLM across several editing timesteps.
Recent research on model editing techniques has primarily focused on pure LLMs [17, 18, 19, 20]. However, the additional visual modality and the interactions between visual and textual modalities make these pure LLM editors less suitable. For example, LLM editors such as those in [17, 18, 21], which are based on locate-then-edit methods, assume that the subject in the query is crucial for model reasoning. These methods perform causal mediation analysis on input text queries containing the subject to identify linear layer weights critical to the LLM’s reasoning. However, in vision-dominated tasks such as Visual Question Answering (VQA), where visual inputs often include substantial relevant information, attribution becomes more challenging. As a result, only limited work has explored how visual representations within VLLMs contribute to response generation and has proposed single-shot editing algorithms [22, 23, 24].
In most LLM applications, single-shot model editing is insufficient to keep the model updated. Thus, the concept of lifelong editing has emerged to address the continuous need for model updates [25, 21, 19, 20]. In lifelong editing scenarios, some LLM editors have demonstrated strong performance. Retrieval-based methods, in particular, avoid directly editing the original model and apply on-the-fly edit retrieval and parameter fusion during inference [25, 19, 20]. This approach offers greater robustness to a growing number of edits compared to editors that rely on permanent parameter modifications. For VLLMs, to the best of our knowledge, there is currently no related research on lifelong editing, as illustrated in Figure 1. Visual modality is quite different from text modality since it typically contains more information and is more noisy. Therefore, the approaches working for LLMs cannot be directly applied to VLLMs.
In this paper, we introduce LiveEdit, a novel framework designed for Lifelong vision language model Editing, which bridges the gap between lifelong LLM editing and VLLMs. In this framework, we design a generative low-rank mixture-of-experts combined with hard and soft routing as a powerful VLLM editor. The two key techniques are outlined as follows:
Generation of Low-Rank Experts: Mixture-of-Experts (MoE) is a technique used to solve complex tasks by combining multiple “experts”, with each expert being responsible for specific data patterns or sub-tasks [26, 27]. In the context of VLLM editing, we treat VLLM’s adherence to a single editing sample as a sub-task. Thus, in lifelong editing, each new incoming editing sample corresponds to a specific low-rank expert, which is responsible for adjusting the model’s response accordingly. Typically, each component in MoE requires training on sub-tasks to obtain its weights. However, directly fitting on editing samples results in poor generalization. Furthermore, performing separate training for each new editing sample is inefficient. Based on the above analysis, we construct an expert generator to generate low-rank experts for each new editing sample. To ensure that the generated experts can effectively refine the VLLM response, we train the expert generator to align the VLLM, with the relevant editing metrics: reliability, textual and modal generality, and textual and modal locality [22]. We store the generated experts in an expert repository. For a new input instance, LiveEdit will select relevant experts from the repository and merge their adapted responses using hard and soft routing, as detailed below.
Hard and Soft Routing: To route and utilize these experts during inference, we propose a two-stage routing strategy. In the first phase, we extract visual features relevant to the edit sample by interacting between the prompt and the visual representation. For each incoming sample, we do the same visual feature extraction and compare it with those of the edit samples to filter out visually irrelevant experts, thereby routing the input instance only to visually relevant experts. This hard routing effectively removes irrelevant visual and textual noise. However, it does not take into account the specific semantics within the edit text, and this coarse-grained routing may lead to the selection of multiple visually matched but text-irrelevant editing samples. In the second stage, to suppress irrelevant samples, we implement multi-expert fusion, or soft routing, by computing the similarity between the input query and the edit text applied in the first stage. By combining absolute and relative weights, relevant and irrelevant editing samples are assigned higher and lower weights, respectively.
In the experiments, the proposed LiveEdit framework is tested with 1, 10, 100, and 1,000 edits on the LLaVA-V1.5 (7B) [28], MiniGPT-4 (7B) [29], and BLIP2-OPT (2.7B) [4] backbones across the E-VQA [22], E-IC [22], and VLKEB [30] benchmark datasets. Comparisons with other strong editors demonstrated the superiority of our approach.
2 Related Works
2.1 Vision Large Language Models
Motivated by recent achievements of LLMs [31], researchers have invested substantial effort in merging LLMs with vision models [32]. VLLMs synchronize pre-trained image encoders, usually a Vision Transformer (ViT) [33], with an LLM decoder. Consequently, this configuration produces a model proficient in handling images alongside text inputs [4, 5, 6, 34, 7]. The training process for VLLMs typically unfolds in a two-phase approach. Initially, an alignment component, which may be a feed forward network [28] or more sophisticated structures such as a resampler [35, 4], is developed to bridge the image encoder with the LLM. This component is trained using pairs of images and their corresponding captions, effectively mapping image tokens onto the input space of the LLM. Subsequently, the focus shifts to broad-spectrum inference capabilities. The model is then refined through exposure to a diverse array of tasks, encompassing visual question-answering scenarios [36, 37] and instruction-based interactions in both visual and textual contexts [38, 39], thereby enriching its functional versatility. However, despite the broad application potential mentioned above, VLLMs still rely on meticulous fine-tuning and editing to ensure adaptability and accuracy across diverse scenarios.
2.2 Model Editing
Model Editing for LLMs: We offer a concise summary of techniques for editing LLMs. (1) Locate-Then-Edit: This method first pinpoints the parameters connected to particular knowledge, then directly modifies these targeted parameters [40]. ROME [17] uses causal mediation analysis to pinpoint the section requiring editing. MEMIT [18] and WILKE [21], building on ROME, facilitate the multiple knowledge editing. (2) Meta Learning: This approach uses a hyper-network to derive the updated weights necessary for editing [41, 42, 43, 44]. (3) Additional Parameters: This technique adds extra trainable parameters to LLMs, training them specifically on a dataset of edits while keeping the original parameters intact [45, 46]. (4) Adding Extra Modules: This strategy memorizes all editing instances and leverages a retrieval mechanism to fetch the most relevant pieces of information for each query [47, 25, 48, 49, 19].
Model Editing for VLLMs: Leveraging multimodal data for knowledge editing on VLLMs better resonates with practical contexts. Previously, the outlined methods were tailored for LLMs, operating solely on single-modal data. Yet, when it comes to knowledge editing on VLLMs, employing multimodal data offers a closer approximation to real-life settings. In the literature, MMEdit [22] and VLKEB [30] contribute novel datasets designed specifically for multimodal knowledge editing tasks. [24] use attribution analysis to explore how VLLMs extract key information from visual representations to generate responses, leading to the design of a single-step VLLM editing technique.
2.3 Mixture-of-Experts (MoE)
The MoE technique [26, 27] decomposes complex tasks into simpler ones, using dedicated models called experts. Recently, MoE layers have been integrated into transformer architectures. For instance, GShard [50] utilized MoE in transformer, achieving significant improvements in machine translation for 100 languages. Switch Transformers [51] further scaled language models with a trillion parameters through efficient MoE designs. However, naive MoE training may cause load imbalance, wherein a few experts are overused while others are underutilized. To combat this, various strategies such as the BASE layer [52], HASH layer [53], and Expert Choice [54] have been developed to optimize MoE models’ capacity. Recent efforts focus on training a decoder-only MoE model with a modified UL2 objective [55]. Notably, Mixtral [56] enhances decision-making by employing token-choice routing to select two out of eight experts, improving overall performance. Our work approaches the MoE structure from a different angle, treating each knowledge update as a mini-task and leveraging low-rank experts to store knowledge for VLLM editing.
Figure 2: Illustration of the LiveEdit framework. The upper part illustrates the editing process of LiveEdit. At time step $t$, the representation of an editing sample $(v_{e_{t}},p_{e_{t}},o_{e_{t}})$ at layer $l_{e}$ serves as an editing signal to generate the editing expert $(U_{e_{t}},V_{e_{t}})$ via $f_{eg}$ and routing features $(\hat{\phi}_{v_{e_{t}}},\hat{\psi}_{p_{e_{t}}})$ via $\hat{f}_{fe}$. Both are then added to the expert repository. The lower part shows the VLLM inference process with LiveEdit, where $\bar{f}_{fe}$ extracts input sample features at layer $l_{e}$ to route editing experts, which then adapt the representation.
3 Preliminaries and Task Definition
We formally define the VLLM editing task and its lifelong extension. Next, we introduce the evaluation criteria.
A VLLM $f_{\theta}:\mathcal{V}\times\mathcal{P}\rightarrow\mathcal{O}$ can be considered a function that maps an image-prompt pair $(v,p)$ to a textual output $o=f_{\theta}(v,p)$. Given an edit sample $(v_{e},p_{e},o_{e})$, where $f_{\theta}(v_{e},p_{e})\neq o_{e}$, a VLLM editor $\text{ME}:\mathcal{F}\times\mathcal{V}\times\mathcal{P}\times\mathcal{O}\rightarrow\mathcal{F}$ produces an updated VLLM $f_{\theta^{\prime}}=\text{ME}(f_{\theta},v_{e},p_{e},o_{e})$. Starting from an initial VLLM $f_{\theta_{0}}$, ME iteratively applies edits as new editing requirements arise in a lifelong context:
$f_{\theta_{t}}=\text{ME}(f_{\theta_{t-1}},v_{e_{t}},p_{e_{t}},o_{e_{t}}),t=1,2,3,...$
At any timestep $t$, an effective ME should ensure that $f_{\theta_{t}}$ satisfies the following three criteria, as outlined in [22]:
Reliability measures the accuracy of the modified model’s responses on edited samples:
$\mathbb{E}_{(v_{e},p_{e},o_{e})\sim\{(v_{e_{\tau}},p_{e_{\tau}},o_{e_{\tau}})\}_{\tau=1}^{t}}\mathbb{I}\{f_{\theta_{t}}(v_{e},p_{e})=o_{e}\}$
where $\mathbb{I}$ is the indicator function that evaluates to 1 when the condition is true.
Generality requires $f_{\theta_{t}}$ can also adapt to relevant variations (e.g., rephrased prompts) in the edited samples, including modal and text generality:
$\mathbb{E}_{(v_{e},p_{e},o_{e})\sim\{(v_{e_{\tau}},p_{e_{\tau}},o_{e_{\tau}})\}_{\tau=1}^{t}}\mathbb{E}_{v_{g}\sim\mathcal{G}(v_{e})}\mathbb{I}\left\{f_{\theta_{t}}(v_{g},p_{e})=o_{e}\right\}$
$\mathbb{E}_{(v_{e},p_{e},o_{e})\sim\{(v_{e_{\tau}},p_{e_{\tau}},o_{e_{\tau}})\}_{\tau=1}^{t}}\mathbb{E}_{p_{g}\sim\mathcal{G}(p_{e})}\mathbb{I}\left\{f_{\theta_{t}}(v_{e},p_{g})=o_{e}\right\}$
where $\mathcal{G}(\cdot)$ represents the relevant neighbors.
Locality requires $f_{\theta_{t}}$ remains consistent with $f_{\theta_{0}}$ for samples unrelated to edits, including modal and text locality:
$\mathbb{E}_{(v_{e},p_{e},o_{e})\sim\{(v_{e_{\tau}},p_{e_{\tau}},o_{e_{\tau}})\}_{\tau=1}^{t}}\mathbb{E}_{(v_{l},p_{l})\sim\mathcal{L}(v_{e},p_{e})}\mathbb{I}_{l}(v_{l},p_{l})$
$\mathbb{E}_{(v_{e},p_{e},o_{e})\sim\{(v_{e_{\tau}},p_{e_{\tau}},o_{e_{\tau}})\}_{\tau=1}^{t}}\mathbb{E}_{p_{l}\sim\mathcal{L}(p_{e})}\mathbb{I}_{l}(\emptyset,p_{l})$
[…]
where $\mathcal{L}(\cdot)$ represents the irrelevant samples.
4 The Proposed LiveEdit Framework
In this section, we formally introduce the LiveEdit framework, with the overall architecture shown in Figure 2. First, we explain how to generate corresponding experts for editing samples and how to extract the semantic features of the editing prompt. We then describe their use in extracting key visual features from the visual representation, thereby maintaining the expert repository. Next, we describe how LiveEdit routes experts during a single inference of the VLLM to instantly adjust its response. Finally, we elaborate on the overall training process of the LiveEdit model.
4.1 Construction and Update of Expert Repository
Since the LLM transformer is the primary module for semantic understanding and response generation, in this work, we consider inserting the MoE editor between the layers of the transformer for editing. Previous work [24] on representation attribution in VLLM editing has shown that the latter layers in the model leverage the prompt semantics to extract relevant visual information to generate responses. Following their findings, we deploy our editor in a high-contribution layer $l_{e}$ within the transformer. The expert repository is initially set to $\mathcal{E}_{0}=\{\}$ and is updated from $\mathcal{E}_{t-1}$ to $\mathcal{E}_{t}$ at timestep $t$ as a new editing sample $(v_{e_{t}},p_{e_{t}},o_{e_{t}})$ is input into the model.
Specifically, given a VLLM $f_{\theta}$, the image and text of the edit sample are converted into embeddings, concatenated, and fed into the transformer. Let $h^{l_{e}}\in\mathbb{R}^{N\times d}$ represent the intermediate output at layer $l_{e}$, where $N$ and $d$ correspond to the sequence length and the intermediate dimension, respectively. Let $h_{v_{e_{t}}}^{l_{e}}$, $h_{p_{e_{t}}}^{l_{e}}$, $h_{o_{e_{t}}}^{l_{e}}$ denote the respective representations of $v_{e_{t}}$, $p_{e_{t}}$, and $o_{e_{t}}$. We define $f_{eg}(\cdot)$ to extract the editing signal and generate the expert:
$(U_{e_{t}},V_{e_{t}})$ $=f_{eg}(h_{v_{e_{t}}}^{l_{e}}\oplus h_{p_{e_{t}}}^{l_{e}}\oplus h_{o_{e_{t}}}^{l_{e}}),$ (1)
[…]
where $U\in\mathbb{R}^{r\times d_{m}}$ and $V\in\mathbb{R}^{r\times d_{m}}$ are two trainable matrices. $\oplus$ denotes concatenation. $r$ and $d_{m}$ are hyper-parameters representing the number of ranks and the module dimension, respectively. The cross attention $\text{CA}(\cdot)$ is formulated as:
$\text{CA}(x,y)=\delta\left(xW_{q}\left(yW_{k}\right)^{T}\right)\cdot yW_{v}$ (2)
where $\delta$ denotes softmax, and $W_{q},W_{k},W_{v}$ are matrices that map inputs into query, key, and value spaces, respectively.
To perform both hard and soft routing for experts, we respectively extract semantics from the prompt representation and obtain key visual information from the visual representation through a feature extractor $\hat{f}_{fe}(\cdot)$:
$(\hat{\phi}_{v_{e_{t}}},\hat{\psi}_{p_{e_{t}}})$ $=\hat{f}_{fe}(h_{v_{e_{t}}}^{l_{e}},h_{p_{e_{t}}}^{l_{e}})$ (3)
[…]
where $\phi\in\mathbb{R}^{1\times kd_{m}}$ and $\psi\in\mathbb{R}^{1\times kd_{m}}$ are trainable feature extraction vectors, and $k$ controls the dimension of the vectors. The extracted features $\hat{\phi}_{v_{e_{t}}}$ and $\hat{\psi}_{p_{e_{t}}}$ have the same shape as $\phi$ and $\psi$. Finally, the expert repository is updated by inserting the group of experts and the routing features as $\mathcal{E}_{t}=\mathcal{E}_{t-1}\cup\{(U_{e_{t}},V_{e_{t}},\hat{\phi}_{v_{e_{t}}},\hat{\psi}_{p_{e_{t}}})\}$.
4.2 Expert Routing and Editing on the Fly
Given an input image-prompt pair $(v_{i},p_{i})$, let its output at the $l_{e}$-th layer be $h^{l_{e}}$, where $h_{v_{i}}^{l_{e}}$ and $h_{p_{i}}^{l_{e}}$ denote the components corresponding to $v_{i}$ and $p_{i}$, respectively. We use an additional feature extraction function $\bar{f}_{fe}$ (defined in consistency within Eq.3, but taking different inputs) to extract routing features from the input:
$(\bar{\phi}_{v_{i}},\bar{\psi}_{p_{i}})=\bar{f}_{fe}(h_{v_{i}}^{l_{e}},h_{p_{i}}^{l_{e}}).$ (4)
Given $\mathcal{E}_{t}=\{(U_{e_{\tau}},V_{e_{\tau}},\hat{\phi}_{v_{e_{\tau}}},\hat{\psi}_{p_{e_{\tau}}})\}_{\tau=1}^{t}$, we filter for experts that are highly relevant to the input sample’s visual content by calculating the similarity between the key visual features of the input and the editing samples:
^=$\hat{\mathcal{E}}=\left\{(U_{e_{\tau}},V_{e_{\tau}},\hat{\psi}_{p_{e_{\tau}}})\mid\bar{\phi}_{v_{i}}\hat{\phi}_{v_{e_{\tau}}}^{T}>\bar{\phi}_{v_{i}}\bar{\phi}_{\Theta}^{T},\;\tau=1,...,t\right\},$ (5)
[…]
where a trainable vision sentinel $\Theta\in\mathbb{R}^{N_{v}\times d}$ is set to dynamically determine the filtering threshold, following [20], which effectively avoids the bias caused by manually set thresholds. $N_{v}$ is the vision token count of $f_{\theta}$. Intuitively, if the input sample is more visually similar to the editing sample than the visual sentinel, then this editing sample should not be selected.
Although the above process effectively selects visually relevant editing experts, some results may still have low prompt semantic relevance. We further use the similarity between prompt features to achieve multi-expert fusion. Thus, the post-edit representation $\hat{h}^{l_{e}}$ is obtained as follows:
$\hat{h}^{l_{e}}=h^{l_{e}}+\!\!\!\!\!\sum\limits_{(U_{e},V_{e},\hat{\psi}_{p_{e}})\in\hat{\mathcal{E}}}\!\!\!\!\!\!\!\!f_{sr}(\bar{\psi}_{p_{i}},\hat{\psi}_{p_{e}},\hat{\mathcal{E}})\rho(h^{l_{e}}U_{e}^{T})V_{e}$ (6)
[…]
where $\rho(\cdot)$ and $\sigma(\cdot)$ are ReLU and sigmoid, respectively. The inner products are rescaled by $\sqrt{d_{m}}$, which is omitted above. The $\hat{h}^{l_{e}}$ will proceed to complete the subsequent layer inference and generate the modified response. The Soft Routing function $f_{sr}(\cdot)$ multiplies absolute weights from the sigmoid and relative weights from the softmax. The absolute weights control the output strength of each expert based on similarity. The relative weights balance the similarity among the selected experts to constrain the scale of the fused residual output within 1, preventing the combined output from generating an excessively large norm.
4.3 Training of LiveEdit
The training primarily consists of two parts: the editing loss, which ensures that the generated MoEs effectively guide the VLLM to follow the editing instructions, and the routing loss, which ensures hard and soft MoE routing. Given a batch of editing samples $\mathcal{D}_{e}=\{(v_{e_{b}},p_{e_{b}},o_{e_{b}})\}_{b=1}^{B}$, and their corresponding sampled generality and locality samples $\mathcal{D}_{g}=\{(v_{g_{b}},p_{g_{b}},o_{g_{b}})\}_{b=1}^{B}$ and $\mathcal{D}_{l}=\{(v_{l_{b}},p_{l_{b}},o_{l_{b}})\}_{b=1}^{B}$, the losses are formulated as follows.
4.3.1 Editing Loss
We mix the experts for the entire batch of editing samples to simulate the scenario during inference, where hard routing leads to multiple experts. First, we obtain the $l_{e}$-th layer outputs $\{(h_{v_{e_{b}}}^{l_{e}},h_{p_{e_{b}}}^{l_{e}},h_{o_{e_{b}}}^{l_{e}})\}_{b=1}^{B}$ for each part of an editing sample in $\mathcal{D}_{e}$. Then, through Eqs. 1 and 3, their corresponding experts and routing features can be obtained as $\{(U_{e_{b}},V_{e_{b}})\}_{b=1}^{B}$ and $\{(\hat{\phi}_{v_{e_{b}}},\hat{\psi}_{p_{e_{b}}})\}_{b=1}^{B}$. We define the expert set for soft routing fusion as ^=$\hat{\mathcal{E}}=\{(U_{e_{b}},V_{e_{b}},\hat{\psi}_{p_{e_{b}}})\}_{b=1}^{B}$. Thus, for any input sample, its representation at layer $l_{e}$ will be modified as in Eq. 6. Defining the VLLM modified in this way as $f_{\theta_{\hat{\mathcal{E}}}}$, the editing loss is defined as follows:
$\ell_{\text{edit}}=\frac{1}{B}\sum\limits_{b=1}^{B}\left(\ell^{(b)}_{\text{rel}}+\ell^{(b)}_{\text{gen}}+\ell_{\text{loc}}^{(b)}\right)$ (8)
where
$\ell^{(b)}_{\text{rel}}=-\log f_{\theta_{\hat{\mathcal{E}}}}\left(o_{e_{b}}\mid v_{e_{b}},p_{e_{b}}\right)$ (9)
$\ell^{(b)}_{\text{gen}}=-\log f_{\theta_{\hat{\mathcal{E}}}}\left(o_{g_{b}}\mid v_{g_{b}},p_{g_{b}}\right)$ (10)
$\ell^{(b)}_{\text{loc}}=\text{KL}\left(f_{\theta}\left(o_{l_{b}}\mid v_{l_{b}},p_{l_{b}}\right)\,||\,f_{\theta_{\hat{\mathcal{E}}}}\left(o_{l_{b}}\mid v_{l_{b}},p_{l_{b}}\right)\right)$ (11)
Here, KL denotes the Kullback-Leibler divergence.
4.3.2 Routing Loss
In the routing part, we maximize the feature similarity between samples within the generality domain, while minimizing the feature similarity between unrelated samples. First, we randomly assign samples within the same generalization domain (i.e., editing samples and their corresponding generalization samples) into two new sets, defined as $\hat{\mathcal{D}}_{g}=\{[\mathcal{D}_{e}^{(b)},\mathcal{D}_{g}^{(b)}]_{\pi_{1}^{(b)}}\}_{b=1}^{B}$, and $\bar{\mathcal{D}}_{g}=\{[\mathcal{D}_{e}^{(b)},\mathcal{D}_{g}^{(b)}]_{\pi_{2}^{(b)}}\}_{b=1}^{B}$. Here, $\pi_{1}^{(b)},\pi_{2}^{(b)}\in[0,1]^{B}$ are the random integer vectors applied across the batch. This approach equalizes the reliability and generality of samples in feature matching with the edited and input samples, enhancing routing robustness. We use $\hat{f}_{fe}$ to extract the routing features $\{(\hat{\phi}_{g_{b}},\hat{\psi}_{g_{b}})\}_{b=1}^{B}$ from $\hat{\mathcal{D}}_{g}$ corresponding to edit end, and use $\bar{f}_{fe}$ to extract the routing features $\{(\bar{\phi}_{g_{b}},\bar{\psi}_{g_{b}})\}_{b=1}^{B}$ and $\{(\bar{\phi}_{l_{e}},\bar{\psi}_{l_{e}})\}_{b=1}^{B}$ from $\bar{\mathcal{D}}_{g}$ and $\mathcal{D}_{l}$ corresponding to input end, respectively. The routing loss is formulated as follows:
$\ell_{\text{route}}=\sum\limits_{b=1}^{B}\ell_{hr}^{(b)}+\ell_{sr_{1}}^{(b)}+\ell_{sr_{2}}^{(b)}$ (12)
Hard Routing loss $\ell_{hr}^{(b)}$ is defined as follows:
$\ell_{hr}^{(b)}=f_{nce}(\bar{\phi}_{g_{b}},\hat{\phi}_{g_{b}},\hat{\Phi}\cup\{\bar{\phi}_{\Theta_{g_{b}}}\})+$ (13)
$\;\;\;f_{nce}(\bar{\phi}_{l_{b}},\bar{\phi}_{\Theta_{l_{b}}},\hat{\Phi}\cup\{\bar{\phi}_{\Theta_{l_{b}}}\})$
where $\hat{\Phi}=\{\hat{\phi}_{g_{b}}\}_{b=1}^{B}$. $\bar{\phi}_{\Theta_{g_{b}}}$ and $\bar{\phi}_{\Theta_{l_{b}}}$ represent the features extracted by the corresponding generality and locality data at the input end (as defined in Eq.5) from the vision sentinel. $f_{nce}$ is the InfoNCE loss [57] formulated as:
$f_{nce}(\alpha,\beta_{+},\{\beta_{j}\}_{j=1}^{n})=-\log\frac{\exp(\alpha\beta_{+}^{T})}{\sum_{j=1}^{n}\exp(\alpha\beta_{j}^{T})}$ (14)
We set the temperature to 1, which is omitted here. The above loss function brings the key visual features of data within the generalization domain closer—even closer than those extracted using the visual sentinel. Meanwhile, it pushes the locality inputs further from the generalization domain, making them relatively closer to the features extracted from the visual sentinel.
Soft Routing includes absolute loss $\ell^{(b)}_{sr_{1}}$ and relative loss $\ell^{(b)}_{sr_{2}}$, defined as follows:
$\ell_{sr_{1}}^{(b)}=-\log\sigma(\bar{\psi}_{g_{b}}\hat{\psi}_{g_{b}}^{T})-\log(1-\sigma(\bar{\psi}_{g_{b}}\hat{\psi}_{\backslash g_{b}}^{T})),$ (15)
$\ell_{sr_{2}}^{(b)}=f_{nce}\left(\bar{\psi}_{g_{b}},\hat{\psi}_{g_{b}},\{\hat{\psi}_{g_{j}}\}_{j=1}^{B}\cup\{\hat{\psi}_{l_{j}}\}_{j=1}^{B}\right),$ (16)
where $\hat{\psi}_{\backslash g_{b}}$ represents a feature randomly selected from $\{\hat{\psi}_{g_{j}}\}_{j=1}^{B}\cup\{\hat{\psi}_{l_{j}}\}_{j=1}^{B}\backslash\{\hat{\psi}_{g_{b}}\}$. Thus, the total training loss is: $\ell_{total}=\ell_{edit}+\ell_{route}$. During training, the parameters of the VLLM, $f_{\theta}$, are frozen. The trainable modules are an experts generation module, $f_{eg}$, and two feature extraction modules, $\hat{f}_{fe}$ and $\bar{f}_{fe}$.
Baseline # Edit Editors E-VQA VLKEB
Rel. T-Gen. M-Gen. T-Loc. M-Loc. Average Rel. T-Gen. M-Gen. T-Loc. M-Loc. Average
LLaVA-V1.5 (7B) 1 FT-L 93.88 87.98 80.25 99.61 94.78 91.30 (±0.42) 94.29 87.00 92.22 91.16 91.37 91.21 (±1.09)
FT-M 87.29 76.11 53.23 100.00 96.95 82.72 (±1.05) 76.31 65.57 59.43 100.00 92.35 78.73 (±0.76)
MEND 91.23 90.05 91.29 91.02 90.22 90.76 (±0.64) 92.13 91.28 90.22 89.19 90.13 90.59 (±1.24)
SERAC 89.33 83.72 84.97 82.05 23.78 72.77 (±0.36) 89.77 89.11 87.92 66.68 14.20 69.54 (±0.83)
TP 35.95 36.12 28.65 93.87 97.61 58.44 (±0.33) 50.77 55.70 51.65 87.93 90.43 67.30 (±0.29)
LTE 94.16 93.57 93.59 94.08 86.26 92.33 (±1.56) 94.42 93.57 93.22 86.84 79.69 89.55 (±1.41)
RECIPE 91.37 86.51 87.73 94.27 88.88 89.75 (±1.13) 92.67 92.35 91.01 89.67 82.85 89.71 (±0.57)
LEMoE 93.60 92.77 89.99 99.28 96.98 94.52 (±1.09) 94.85 93.09 91.67 87.03 87.88 90.90 (±0.29)
LiveEdit 94.28 94.51 88.01 100.00 100.00 95.36 (±0.57) 96.43 95.22 93.72 100.00 100.00 97.08 (±0.62)
10 FT-L 90.57 84.14 73.21 95.56 81.50 85.00 (±1.07) 88.05 85.32 85.23 74.53 85.74 83.77 (±1.22)
FT-M 84.90 73.53 49.99 100.00 55.98 72.88 (±0.63) 68.63 57.57 56.56 100.00 82.99 73.15 (±0.23)
MEND 3.58 3.55 3.53 2.10 1.26 2.80 (±0.02) 0.18 0.24 0.05 0.03 0.19 0.14 (±0.00)
SERAC 88.09 83.40 83.57 64.91 15.50 67.10 (±0.92) 81.55 74.49 80.24 54.71 13.15 60.83 (±0.98)
TP 32.71 31.23 28.58 75.10 91.17 51.76 (±0.60) 44.56 47.52 45.36 52.21 66.61 51.25 (±0.69)
LTE 92.83 91.41 90.82 86.38 85.52 89.39 (±0.34) 90.06 81.52 88.11 83.40 81.48 84.91 (±0.78)
RECIPE 90.22 85.92 86.24 90.34 88.11 88.17 (±1.48) 83.92 76.23 82.84 86.33 83.69 82.60 (±0.72)
LEMoE 91.95 86.54 79.82 85.19 49.81 78.66 (±1.03) 91.55 84.58 81.03 67.19 72.81 79.43 (±0.52)
LiveEdit 93.79 93.21 86.42 100.00 100.00 94.68 (±1.03) 95.54 94.52 91.25 100.00 100.00 96.26 (±0.33)
1000 FT-L 71.39 59.83 57.41 55.55 48.99 58.63 (±0.17) 68.14 66.38 66.98 65.61 75.35 68.49 (±0.32)
FT-M 69.57 56.34 44.07 100.00 41.47 62.29 (±0.40) 53.41 48.80 43.16 100.00 57.03 60.48 (±0.50)
MEND 0.04 0.05 0.05 0.08 0.09 0.06 (±0.00) 0.03 0.05 0.07 0.06 0.08 0.06 (±0.00)
SERAC 85.57 75.58 82.01 62.46 15.69 64.26 (±0.37) 60.93 56.49 60.06 52.94 15.04 49.09 (±0.36)
TP 16.56 16.80 15.65 7.28 15.60 14.38 (±0.14) 5.46 4.81 5.51 2.77 7.19 5.15 (±0.07)
LTE 83.93 82.55 81.34 83.97 73.09 80.98 (±1.36) 64.51 56.26 64.80 80.85 76.52 68.59 (±0.60)
RECIPE 87.00 76.81 83.09 86.95 87.03 84.18 (±0.80) 62.00 56.84 61.50 85.37 82.07 69.56 (±0.31)
LEMoE 30.80 25.75 24.32 71.45 46.23 39.71 (±0.23) 67.97 61.07 58.16 48.48 44.06 55.95 (±0.36)
LiveEdit 92.93 90.16 84.30 100.00 96.43 92.76 (±0.20) 92.22 83.97 82.75 100.00 100.00 91.79 (±0.55)
BLIP2-OPT (2.7B) 1 FT-L 52.86 48.80 32.94 98.24 94.27 65.42 (±0.69) 54.31 54.27 54.08 98.40 94.37 71.09 (±1.05)
FT-M 91.70 87.24 33.30 100.00 85.22 79.49 (±0.72) 92.64 80.97 63.62 100.00 83.02 84.05 (±0.70)
MEND 93.13 92.76 93.07 92.00 75.81 89.35 (±0.93) 94.91 93.81 93.84 94.98 86.54 92.82 (±0.82)
SERAC 88.39 84.50 84.25 85.82 26.00 73.79 (±1.01) 87.95 84.67 85.20 68.10 17.75 68.73 (±0.97)
TP 70.14 65.80 53.05 98.11 85.33 74.49 (±0.38) 50.98 49.47 50.88 94.76 78.57 64.93 (±1.02)
LTE 95.74 93.86 86.90 97.93 87.97 92.48 (±0.70) 94.13 91.93 92.23 93.89 92.27 92.89 (±1.01)
RECIPE 89.42 86.24 87.53 99.87 89.16 90.45 (±1.46) 92.38 89.74 89.17 97.13 94.46 92.58 (±1.16)
LEMoE 93.56 92.23 91.40 98.50 85.21 92.18 (±0.73) 94.59 93.14 92.37 94.53 61.53 87.23 (±0.34)
LiveEdit 96.67 94.20 93.82 100.00 100.00 96.94 (±1.32) 98.77 98.08 94.89 100.00 100.00 98.35 (±1.58)
1000 FT-L 45.10 34.62 35.42 48.42 41.24 40.96 (±0.29) 55.39 54.34 53.87 50.80 54.00 53.68 (±0.80)
FT-M 40.40 31.46 27.85 100.00 27.44 45.43 (±0.68) 47.03 49.68 46.99 100.00 41.41 57.02 (±0.13)
MEND 15.84 14.35 17.73 91.74 70.17 41.97 (±0.12) 37.22 38.03 37.19 91.49 84.10 57.61 (±0.58)
SERAC 83.35 70.80 80.32 67.66 13.13 63.05 (±0.87) 53.58 45.78 52.42 56.81 16.90 45.10 (±0.38)
TP 20.63 15.09 18.41 8.65 8.25 14.21 (±0.18) 24.36 24.21 24.25 16.37 19.96 21.83 (±0.14)
LTE 89.32 82.82 81.51 94.86 69.83 83.67 (±1.05) 61.67 51.05 61.60 94.78 90.94 72.01 (±0.66)
RECIPE 84.99 74.20 82.04 96.82 87.73 85.16 (±1.32) 54.64 46.54 54.10 94.60 96.93 69.37 (±1.04)
LEMoE 19.73 17.34 18.22 72.01 31.06 31.67 (±0.14) 34.74 33.43 32.05 55.55 50.04 41.16 (±0.58)
LiveEdit 94.42 91.98 84.65 100.00 97.38 93.69 (±0.67) 97.00 91.92 87.53 100.00 100.00 95.29 (±1.48)
Table 1: Partial results of lifelong edit performance for BLIP2-OPT and LLaVA-V1.5 on the E-VQA and VLKEB datasets. Due to space limitations, please refer to Appendix 8 for the complete results, including those for the E-IC dataset and the MiniGPT-4 model. “Rel.”, “T/M-Gen.” and “T/M-Loc.” stand for reliability, text/modal generality, and text/modal locality, respectively. “# Edit” indicates the number of edits. The t-tests demonstrate our improvements are statistically significant with $p<0.05$ level.
5 Experiments
5.1 Experimental Settings
Datasets: Following [22], we use E-VQA (Editing Visual Question Answering) and E-IC (Editing Image Caption) as evaluation datasets. Additionally, we incorporate VLKEB [30], which is composed of real images to better represent real-world scenarios.
VLLM Backbones: For comprehensive evaluation, we select VLLM backbones based on both model architecture and parameter scale, including BLIP2-OPT (2.7B) [4], LLaVA-V1.5 (7B) [28], and MiniGPT-4 (7B) [29].
Baseline Editors: To our knowledge, there are currently no editors specifically designed for lifelong VLLM editing. Therefore, following [22], in addition to the basic FT-L and FT-M, which respectively fine-tune the final layer of the LLM and the visual encoder, we adapt LLM-based editing techniques to VLLM. These include MEND [42], TP [45], LTE [19], RECIPE [20], and LEMoE [46].
For details on the experimental setup, model hyper-parameters, and training specifics, please refer to Appendix 7. Building on the experimental settings above, we conduct a comprehensive evaluation of edit performance and perform an in-depth analysis of LiveEdit’s internals.
5.2 General Performance of Lifelong Editing
Table 1 shows partial results of lifelong editing experiments. In single-edit scenarios, our method generally achieves optimal performance. Methods such as FT-L, FT-M, and MEND [42] try to modify original model parameters and perform well initially. However, their performance deteriorates with more edits due to overfitting and cumulative parameter scaling [21, 58]. TP [45] addresses parameter scaling by introducing additional neurons, but a single neuron can’t encapsulate each edit’s visual information. Retrieval-based methods (SERAC [47], LTE [19], RECIPE [20]) remain robust in lifelong editing by decoupling edits from model parameters, yet struggle with semantically similar but visually different edit samples. LEMoE [46], a MoE-based editor, excels with few edits but suffers from issues like disruptive greedy routing and limited generalization due to overfitting experts to batch edits. Our method, LiveEdit, surpasses in edit performance. Contrastive learning-based expert routing resolves LEMoE’s issues, enhancing both edit speed and generalization. By decoupling edit samples into independent experts and applying a fusion strategy for vision-related experts, we prevent the semantic conflicts LEMoE encounters from its sequential batching. Importantly, our method maintains nearly 100% locality performance even with increasing edits. Three main factors contribute to our performance: First, hard routing filters out visually irrelevant experts. Second, soft routing scales influence by assigning lower scales to textually irrelevant experts. Finally, the locality editing loss further confines experts’ influence on response adaptation.
Figure 3: The impact of the combination of module dimension $d_{m}$ and expert rank $r$ on LiveEdit’s edit performance. Experiments are conducted on the E-VQA dataset, with 1,000 edits on BLIP2. Circle size represents LiveEdit’s training parameters, and color intensity indicates the average edit performance across five metrics.
5.3 Hyper-Parameter Search
We conducted a comprehensive hyper-parameters search for LiveEdit to select the most suitable combination. The most important hyper-parameters are discussed below.
Trade-off Between Model Scale and Edit Performance: Figure 3 reports the effects of different combinations of module dimension $d_{m}$ and expert rank $r$ on edit performance and model scale. The $d_{m}$ has a significant impact on the parameter count of LiveEdit. In terms of edit performance, both $d_{m}$ and $r$ have substantial effects. Increasing one while keeping the other fixed improves edit performance, though the improvement gradually becomes flat. Additionally, combinations along the diagonal show generally consistent performance. Regarding model scale, $d_{m}$ predominantly influences variation, while $r$ has minimal impact. Based on this analysis, the optimal configuration strategy is to select an appropriate $d_{m}$ and maximize $r$ as much as possible. However, since $r$ linearly controls the growth rate of the expert repository, specific choices should also consider the memory needed to store the expert repository. Additionally, we expand the dimension control parameter $k$ for feature extraction, as shown in Figure 4. Increasing $k$ enhances the feature extraction capability, but an excessively high $k$ may introduce noise, leading to incorrect matches, such as reduced modal locality.
The Attached Layer of LiveEdit: Figure 5 shows the impact of the layer attached by LiveEdit on edit performance. It can be seen that as the layer depth increases, edit performance improves, reaching a peak at 21 layers. After this point, edit performance slightly declines. We speculate that transformers typically perform semantic understanding in the early layers and response generation in the later layers [59, 17, 24]. Attaching to an earlier layer prevents LiveEdit from leveraging VLLM’s semantic understanding capabilities to enhance the feature extraction process. For too deep layers, VLLM has largely stabilized the predictive tendencies of the response, making it more challenging for LiveEdit to adapt. The above results also align with the attribution conclusions of [24].
Figure 4: The dimension control parameter $k$ for feature extraction.
Figure 5: Impact of LiveEdit attached layer index $l_{e}$. Results of 1,000 edits for BLIP2 on E-VQA dataset are reported.
Settings Rel. T-Gen. M-Gen. T-Loc. M-Loc. Average
N/A 20.57 19.03 14.17 100.00 100.00 50.75
LiveEdit 94.42 91.98 84.65 100.00 97.38 93.69
- $\ell_{sr1}$ 87.93 83.77 73.40 100.00 76.93 84.41
- $\ell_{sr2}$ 89.21 85.46 77.34 100.00 91.69 88.74
- SR 88.92 81.49 70.94 100.00 75.66 83.40
HR* 93.60 88.50 80.37 100.00 84.77 89.45
Table 2: Ablation study of LiveEdit.
5.4 Ablation Study
Table 2 presents the ablation results for LiveEdit, where 1000 edits are applied to BLIP2 on the E-VQA dataset. The removal of soft routing (- SR, which directly averages experts) results in a significant performance drop for LiveEdit. We can also observe that a large portion of this loss originates from the removal of the absolute soft routing loss $\ell_{sr_{1}}$, particularly impacting modal locality. This is because all hard-routed experts are assigned a weight to adapt the representation, even if they are irrelevant to the input prompt, ensuring that the sum of adaptation strengths equals 1. Similarly, removing the relative soft routing loss $\ell_{sr_{2}}$ also leads to performance degradation, as the combined absolute weights of multiple experts may exceed 1 during testing, resulting in excessively high values. HR* modifies the visual extraction that leverages prompt semantics by directly compressing the entire visual representation for hard routing. This introduces additional visual noise, leading to the selection of more irrelevant experts, which in turn reduces the efficacy of subsequent steps.
Figure 6: Instance analysis. The left side shows the model input to LLaVa after editing with LiveEdit. The right side displays the top 3 experts based on hard routing. The heatmaps represent the visual regions focused on when extracting key visual features. The bar chart below shows fusion weights from soft routing.
5.5 Instance Analysis
We conducted an instance analysis of LiveEdit, as shown in Figure 6. We perform 100 edits on LLaVa, including samples partially related in visual context to the incoming input samples. The figure reports the top 3 hard routing results for two inputs, as well as the fusion weights these experts received in soft routing. Using a perturbation-based attribution tool [60, 24], we visualize the visual regions focused on for key visual feature extraction. It can be observed that the highlighted regions are closely aligned with the prompt semantics, which filters out irrelevant visual noise and benefits hard routing. The bar chart indicates that experts unrelated to the input query receive very low fusion weights, significantly benefiting the locality of edits.
6 Conclusion
In conclusion, we introduce LiveEdit to bridge the gap between LLM and VLLM editing, with an editing expert generator and a combination of hard/soft routers. The framework successfully addresses the limitation of exiting editors in the VLLM lifelong editing scenarios. Our benchmark and extensive experiments confirm its superiority and highlight the effectiveness of each component, which support more accurate and adaptable VLLM editing in real-world applications.
Supplementary Material
7 Details of Experimental Settings
7.1 Datasets:
E-VQA [22] is a dataset designed to facilitate the correction of error-prone samples in VQA-v2 [36] for VLLMs, containing 6,345 training samples and 2,093 testing samples. The VQA task involves providing a VLLM with an image and a related question, requiring it to analyze both the visual content and the question to produce an accurate textual response.
E-IC [22] is a dataset created for editing VLLMs to rectify errors in COCO Caption [61], consisting of 2,849 training samples and 1,000 testing samples. The IC task requires the model to generate an accurate textual description for a given image, demanding a comprehensive understanding and articulation of the visual content.
Each sample in these two datasets includes an edit sample, along with two additional samples each for modal and text generality, and two samples each for modal and text locality. Generality samples are created by rephrasing images and queries using Stable Diffusion [62] and ChatGLM [63], respectively. Meanwhile, locality samples are generated using unrelated images and queries from the OK-VQA [64] and NQ [42] datasets.
VLKEB [30] is an editor evaluation dataset constructed based on MMKG (Multi-Modal Knowledge Graph) [65]. For each sample, it also includes an editing sample, along with two samples for modal and text generality, and two samples for modal and text locality. It contains 5,000 training samples and 3,174 testing samples.
Unlike the aforementioned datasets, VLKEB uses CLIP [66] to retrieve semantically similar real-world images, which are subsequently verified manually. This process ensures higher image quality compared to datasets relying on image generation models, making VLKEB more representative of real-world scenarios.
7.2 VLLM Backbones:
Below, we introduce the VLLM backbones used in our experiments. For the specific versions utilized, please refer to the corresponding footnotes.
LLaVA [^2] https://huggingface.co/liuhaotian/llava-v1.5-7b [28] is a Vision-Language Large Model (VLLM) that bridges the gap between visual and language understanding by leveraging GPT-4 [67] to create an instruction-tuning dataset for vision-language pre-training. It achieves the transformation of visual features into the textual representation space through a simple yet effective two-layer MLP placed between the visual encoder and the language model LLaMA [1]. This design allows LLaVA to excel at tasks such as visual question answering and instruction following with high efficiency and precision.
BLIP2 [4] introduces a visual query transformer, called Q-Former, which is trained through a two-stage pre-training process to compress visual features into fixed-length representations within the language space. This design bridges the representational gap between the frozen visual encoder [66] and the frozen language model. BLIP2 comes in several variants, and in this paper, we follow [22] to experiment with BLIP2-OPT [^3] https://huggingface.co/Salesforce/blip2-opt-2.7b. While LLaVa utilizes the full spectrum of visual features and preserves more fine-grained details, BLIP2 significantly reduces the length of visual representations, thereby improving the inference efficiency.
MiniGPT-$4$https://huggingface.co/Vision-CAIR/MiniGPT-4 [29], developed based on BLIP2, keeps the Q-Former and the language model Vicuna [38] frozen, training only a linear layer positioned after the Q-Former to efficiently align visual features with Vicuna.
Editors Backbones Edit Iterations Optimizer Learning Rate Edit Layers
FT-L LLaVA-V1.5 (7B) 25 AdamW 1e-3 The last layer of the language transformer
BLIP2-OPT (2.7B) 25 AdamW 1e-3 The last layer of the language transformer
MiniGPT-4 (7B) 25 AdamW 1e-3 The last layer of the language transformer
FT-M LLaVA-V1.5 (7B) 25 AdamW 1e-3 The multi modal projector
BLIP2-OPT (2.7B) 25 AdamW 1e-3 Qformer
MiniGPT-4 (7B) 25 AdamW 1e-3 Qformer
TP LLaVA-V1.5 (7B) 25 Adam 1e-2 The last layer of the language transformer
BLIP2-OPT (2.7B) 25 Adam 1e-2 The last layer of the language transformer
MiniGPT-4 (7B) 25 Adam 1e-2 The last layer of the language transformer
LEMoE LLaVA-V1.5 (7B) 100 AdamW 2e-4 The 18-th layer of the language transformer
BLIP2-OPT (2.7B) 100 AdamW 2e-4 The 18-th layer of the language transformer
MiniGPT-4 (7B) 100 AdamW 2e-4 The 18-th layer of the language transformer
Table 3: Editing details of direct editing editors.
Editors Backbones Optimizer Learning Rate Edit/Modified Layers
MEND LLaVA-V1.5 (7B) Adam 1e-4 layer 29,30,31 of the language transformer
BLIP2-OPT (2.7B) Adam 1e-4 layer 29,30,31 of the language transformer
MiniGPT-4 (7B) Adam 1e-4 layer 29,30,31 of the language transformer
SERAC LLaVA-V1.5 (7B) Adam 1e-4 -
BLIP2-OPT (2.7B) Adam 1e-4 -
MiniGPT-4 (7B) Adam 1e-4 -
LTE LLaVA-V1.5 (7B) Adam 5e-6 The whole language transformer
BLIP2-OPT (2.7B) Adam 5e-6 The whole language transformer
MiniGPT-4 (7B) Adam 5e-6 The whole language transformer
RECIPE LLaVA-V1.5 (7B) Adam 1e-5 -
BLIP2-OPT (2.7B) Adam 1e-5 -
MiniGPT-4 (7B) Adam 1e-5 -
Table 4: Editing details of training-based editing methods.
7.3 Baseline Editors:
FT (Fine-Tuning) comprises two variants: FT-L and FT-M. According to [22], FT-L involves fine-tuning the final layer of the language transformer within the VLLM, whereas FT-M focuses on fine-tuning the visual encoder module of the VLLM for a given edit sample.
MEND (Model Editor Networks with Decomposition) [42] trains a set of small MLP hyper-networks to achieve efficient model editing. These hyper-networks generate the FFN matrix parameter offsets by inputting the decomposed back-propagation gradients of these matrices on edit samples. Through editing-specific training, these hyper-networks can produce matrix offsets that enable LLMs to satisfy editing requirements.
SERAC (Semi-parametric Editing with a Retrieval-Augmented Counterfactual) [47] is a memory-based method. It trains a scope classifier and a small counterfactual model while storing edit samples in memory. The classifier determines whether subsequent inputs are related to the edit samples. If they are related, the inputs are sent to the counterfactual model for response modification; otherwise, they are passed to the original model to generate responses. In our experiments, we follow the setup of MMEdit [22], where BERT [^5] https://huggingface.co/google-bert/bert-base-cased [68] and OPT-125M [^6] https://huggingface.co/facebook/opt-125m [69] are separately trained as the scope classifier and the counterfactual model.
TP (Transformer Patcher) [45] assumes that each neuron in the LLM can carry a piece of knowledge. Therefore, for every new piece of edit sample, TP inserts and trains a new neuron in the final layer of the FFN within the LLM to accommodate the edit sample.
LTE (Learning To Edit) [19] performs supervised fine-tuning of LLMs on edit samples, enabling the model to adapt its response when an edit sample is provided as a contextual prefix. To address sequential editing scenarios, it stores edit samples in a retrieval database and uses a pre-existing text retriever [70] to retrieve edit samples relevant to the input.
RECIPE (REtrieval-augmented ContInue Prompt lEarning) [20] aims to explore short editing prefixes to achieve lifelong editing for LLMs. By training an editing prompt generator, edit samples are converted into short continuous prompts, which serve as prefix contexts for LLMs to adapt their responses. Additionally, RECIPE trains an editing retriever to retrieve relevant edited samples based on the input.
LEMoE (Lifelong Editing with Mixture of Experts) [46] achieves lifelong editing for LLMs by inserting and maintaining a MoE in the final layer of the LLM. In lifelong editing, whenever a batch of editing requirements arises, an editing expert is initialized and trained to adapt the model’s responses for that specific batch of edit samples. During the training of a new editing expert, previously trained experts also participate in the back-propagation process, but their parameters remain frozen. Additionally, a key vector is trained alongside the expert to route the input to the relevant experts.
7.4 Model Settings and Training Details
[…]
Other Baselines: The hyperparameters of TP [45], LTE [19], and LEMoE [46] are adopted from their respective papers. They are slightly adjusted to accommodate VLLM editing, such as mapping visual representations to the language representation space via a trainable MLP module. For the other baselines, we follow [22]. For the experimental details of direct editing methods, please refer to Table 3. For training-based editing methods, we also set the maximum number of iterations to 200K. Training details are provided in Table 4.
8 The Complete Lifelong Editing Results
Tables 5 and 6 present the lifelong editing performance of various editors on the E-VQA [22], VLKEB [22], and E-IC [22] datasets, with LLaVA [28], BLIP2 [4], and MiniGPT-4 [29] as backbones, respectively. The results demonstrate a similar overall trend, confirming the effectiveness of our method. It is worth noting, however, that all methods exhibit relatively lower performance on the E-IC dataset. Since the image captioning task typically requires adapting descriptions of the entire image, editing involves carrying more complex information. This leads to a decline in the editing efficacy of the editors due to the limitations in representational power.
Baseline # Edit Editors E-VQA VLKEB E-IC
Rel. T-Gen. M-Gen. T-Loc. M-Loc. Average Rel. T-Gen. M-Gen. T-Loc. M-Loc. Average Rel. T-Gen. M-Gen. T-Loc. M-Loc. Average
LLaVA (7B) 1 FT-L 93.88 87.98 80.25 99.61 94.78 91.30 (±0.42) 94.29 87.00 92.22 91.16 91.37 91.21 (±1.09) 73.48 72.98 65.79 99.28 99.06 82.12 (±0.82)
FT-M 87.29 76.11 53.23 100.00 96.95 82.72 (±1.05) 76.31 65.57 59.43 100.00 92.35 78.73 (±0.76) 56.19 56.55 49.94 100.00 100.00 72.54 (±0.85)
MEND 91.23 90.05 91.29 91.02 90.22 90.76 (±0.64) 92.13 91.28 90.22 89.19 90.13 90.59 (±1.24) 92.82 91.81 90.59 96.38 93.69 93.06 (±1.40)
SERAC 89.33 83.72 84.97 82.05 23.78 72.77 (±0.36) 89.77 89.11 87.92 66.68 14.20 69.54 (±0.83) 88.18 81.03 85.61 84.01 28.58 73.48 (±1.19)
TP 35.95 36.12 28.65 93.87 97.61 58.44 (±0.33) 50.77 55.70 51.65 87.93 90.43 67.30 (±0.29) 57.63 59.23 55.34 60.90 88.00 64.22 (±0.49)
LTE 94.16 93.57 93.59 94.08 86.26 92.33 (±1.56) 94.42 93.57 93.22 86.84 79.69 89.55 (±1.41) 93.35 91.30 92.77 95.77 91.98 93.03 (±0.55)
RECIPE 91.37 86.51 87.73 94.27 88.88 89.75 (±1.13) 92.67 92.35 91.01 89.67 82.85 89.71 (±0.57) 84.45 76.97 81.57 96.53 96.37 87.18 (±1.08)
LEMoE 93.60 92.77 89.99 99.28 96.98 94.52 (±1.09) 94.85 93.09 91.67 87.03 87.88 90.90 (±0.29) 93.80 91.42 90.61 95.14 93.00 92.79 (±0.41)
LiveEdit 94.28 94.51 88.01 100.00 100.00 95.36 (±0.57) 96.43 95.22 93.72 100.00 100.00 97.08 (±0.62) 82.16 81.01 78.27 100.00 100.00 88.29 (±1.42)
10 FT-L 90.57 84.14 73.21 95.56 81.50 85.00 (±1.07) 88.05 85.32 85.23 74.53 85.74 83.77 (±1.22) 68.74 67.05 60.95 97.05 91.20 77.00 (±0.80)
FT-M 84.90 73.53 49.99 100.00 55.98 72.88 (±0.63) 68.63 57.57 56.56 100.00 82.99 73.15 (±0.23) 57.23 55.92 50.09 100.00 90.31 70.71 (±1.18)
MEND 3.58 3.55 3.53 2.10 1.26 2.80 (±0.02) 0.18 0.24 0.05 0.03 0.19 0.14 (±0.00) 66.49 64.77 58.42 87.25 86.15 72.62 (±1.13)
SERAC 88.09 83.40 83.57 64.91 15.50 67.10 (±0.92) 81.55 74.49 80.24 54.71 13.15 60.83 (±0.98) 56.57 56.88 52.62 59.96 14.66 48.14 (±0.40)
TP 32.71 31.23 28.58 75.10 91.17 51.76 (±0.60) 44.56 47.52 45.36 52.21 66.61 51.25 (±0.69) 45.28 47.25 42.78 19.74 59.59 42.93 (±0.38)
LTE 92.83 91.41 90.82 86.38 85.52 89.39 (±0.34) 90.06 81.52 88.11 83.40 81.48 84.91 (±0.78) 52.16 55.70 48.39 90.74 89.09 67.21 (±0.89)
RECIPE 90.22 85.92 86.24 90.34 88.11 88.17 (±1.48) 83.92 76.23 82.84 86.33 83.69 82.60 (±0.72) 56.00 56.19 52.14 91.80 95.31 70.29 (±0.98)
LEMoE 91.95 86.54 79.82 85.19 49.81 78.66 (±1.03) 91.55 84.58 81.03 67.19 72.81 79.43 (±0.52) 89.00 85.23 83.24 86.39 82.86 85.34 (±0.91)
LiveEdit 93.79 93.21 86.42 100.00 100.00 94.68 (±1.03) 95.54 94.52 91.25 100.00 100.00 96.26 (±0.33) 81.93 80.80 75.55 100.00 100.00 87.66 (±0.49)
100 FT-L 79.67 70.05 64.07 83.47 54.44 70.34 (±0.95) 75.41 73.67 74.16 70.01 82.05 75.06 (±1.17) 65.08 60.90 58.46 86.82 89.19 72.09 (±0.65)
FT-M 82.90 72.83 47.26 100.00 43.39 69.28 (±0.43) 60.60 59.79 56.29 100.00 68.07 68.95 (±0.78) 59.41 55.44 52.16 100.00 72.93 67.99 (±0.43)
MEND 2.22 2.20 2.21 0.21 0.62 1.49 (±0.02) 0.56 0.58 0.66 0.18 0.07 0.41 (±0.00) 56.83 56.89 53.07 87.63 84.64 67.81 (±0.93)
SERAC 88.08 81.53 82.48 62.13 12.90 65.42 (±0.32) 72.25 62.43 70.68 53.73 13.69 54.56 (±0.49) 53.35 53.70 49.41 48.04 17.25 44.35 (±0.50)
TP 29.37 28.72 24.66 14.64 45.01 28.48 (±0.33) 19.71 20.07 19.36 11.40 24.05 18.92 (±0.08) 22.88 25.81 20.90 3.59 14.87 17.61 (±0.06)
LTE 88.92 87.89 85.72 84.34 81.60 85.69 (±0.37) 80.27 64.25 80.13 81.62 79.11 77.08 (±0.85) 48.37 49.63 45.08 87.66 87.62 63.67 (±0.60)
RECIPE 89.86 83.32 84.82 87.37 85.08 86.09 (±0.55) 73.97 63.73 72.69 86.20 82.59 75.84 (±0.58) 53.23 53.75 49.36 87.64 95.52 67.90 (±0.36)
LEMoE 42.41 36.60 34.33 78.57 53.28 49.04 (±0.45) 83.07 75.36 71.72 54.09 49.68 66.78 (±0.60) 55.36 53.14 51.12 87.77 81.35 65.75 (±1.02)
LiveEdit 93.54 92.34 85.89 100.00 99.31 94.21 (±0.34) 94.56 90.65 89.56 100.00 100.00 94.95 (±1.34) 80.81 78.77 63.52 100.00 100.00 84.62 (±0.62)
1000 FT-L 71.39 59.83 57.41 55.55 48.99 58.63 (±0.17) 68.14 66.38 66.98 65.61 75.35 68.49 (±0.32) 59.78 54.99 54.17 65.37 78.96 62.65 (±0.43)
FT-M 69.57 56.34 44.07 100.00 41.47 62.29 (±0.40) 53.41 48.80 43.16 100.00 57.03 60.48 (±0.50) 49.21 47.75 43.81 100.00 35.14 55.18 (±0.88)
MEND 0.04 0.05 0.05 0.08 0.09 0.06 (±0.00) 0.03 0.05 0.07 0.06 0.08 0.06 (±0.00) 54.39 54.14 50.99 83.87 80.60 64.80 (±0.35)
SERAC 85.57 75.58 82.01 62.46 15.69 64.26 (±0.37) 60.93 56.49 60.06 52.94 15.04 49.09 (±0.36) 52.93 53.44 49.01 49.91 16.65 44.39 (±0.73)
TP 16.56 16.80 15.65 7.28 15.60 14.38 (±0.14) 5.46 4.81 5.51 2.77 7.19 5.15 (±0.07) 10.28 13.14 9.75 1.71 4.45 7.87 (±0.13)
LTE 83.93 82.55 81.34 83.97 73.09 80.98 (±1.36) 64.51 56.26 64.80 80.85 76.52 68.59 (±0.60) 48.83 49.96 45.68 85.17 86.41 63.21 (±0.67)
RECIPE 87.00 76.81 83.09 86.95 87.03 84.18 (±0.80) 62.00 56.84 61.50 85.37 82.07 69.56 (±0.31) 53.11 53.48 48.99 87.93 94.84 67.67 (±1.11)
LEMoE 30.80 25.75 24.32 71.45 46.23 39.71 (±0.23) 67.97 61.07 58.16 48.48 44.06 55.95 (±0.36) 34.50 31.38 28.14 82.09 75.88 50.40 (±0.26)
LiveEdit 92.93 90.16 84.30 100.00 96.43 92.76 (±0.20) 92.22 83.97 82.75 100.00 100.00 91.79 (±0.55) 72.80 69.95 57.05 100.00 99.79 79.92 (±0.72)
BLIP2 (2.7B) 1 FT-L 52.86 48.80 32.94 98.24 94.27 65.42 (±0.69) 54.31 54.27 54.08 98.40 94.37 71.09 (±1.05) 45.02 44.47 40.72 99.02 98.27 65.50 (±0.49)
FT-M 91.70 87.24 33.30 100.00 85.22 79.49 (±0.72) 92.64 80.97 63.62 100.00 83.02 84.05 (±0.70) 67.14 61.76 43.34 100.00 96.76 73.80 (±0.16)
MEND 93.13 92.76 93.07 92.00 75.81 89.35 (±0.93) 94.91 93.81 93.84 94.98 86.54 92.82 (±0.82) 94.96 92.45 92.33 94.95 88.86 92.71 (±1.45)
SERAC 88.39 84.50 84.25 85.82 26.00 73.79 (±1.01) 87.95 84.67 85.20 68.10 17.75 68.73 (±0.97) 88.71 83.81 84.38 84.28 24.70 73.18 (±0.25)
TP 70.14 65.80 53.05 98.11 85.33 74.49 (±0.38) 50.98 49.47 50.88 94.76 78.57 64.93 (±1.02) 49.65 48.58 46.02 93.69 78.95 63.38 (±1.00)
LTE 95.74 93.86 86.90 97.93 87.97 92.48 (±0.70) 94.13 91.93 92.23 93.89 92.27 92.89 (±1.01) 92.58 91.94 90.90 97.80 91.42 92.93 (±1.37)
RECIPE 89.42 86.24 87.53 99.87 89.16 90.45 (±1.46) 92.38 89.74 89.17 97.13 94.46 92.58 (±1.16) 85.20 81.44 82.71 100.00 94.59 88.79 (±1.23)
LEMoE 93.56 92.23 91.40 98.50 85.21 92.18 (±0.73) 94.59 93.14 92.37 94.53 61.53 87.23 (±0.34) 93.07 91.37 83.28 94.45 60.44 84.52 (±0.38)
LiveEdit 96.67 94.20 93.82 100.00 100.00 96.94 (±1.32) 98.77 98.08 94.89 100.00 100.00 98.35 (±1.58) 80.60 80.12 76.88 100.00 100.00 87.52 (±0.31)
10 FT-L 53.77 48.91 37.66 97.68 80.40 63.68 (±0.88) 55.81 55.55 55.29 93.37 75.45 67.09 (±0.71) 46.20 44.07 40.54 98.37 92.59 64.35 (±0.75)
FT-M 85.27 81.09 34.60 100.00 49.00 69.99 (±0.15) 92.48 86.02 71.48 100.00 60.34 82.06 (±0.47) 63.79 56.88 47.75 100.00 52.41 64.17 (±0.42)
MEND 40.40 32.21 27.73 87.92 67.74 51.20 (±0.39) 54.47 52.59 53.61 91.96 84.58 67.44 (±0.16) 7.15 7.48 7.01 19.74 21.70 12.62 (±0.05)
SERAC 86.12 82.24 82.52 70.24 16.06 67.43 (±1.10) 78.43 67.72 75.12 56.21 14.54 58.40 (±0.73) 47.69 46.94 43.46 65.29 16.65 44.01 (±0.15)
TP 55.86 52.39 43.00 91.26 63.62 61.22 (±0.54) 50.19 51.61 50.13 89.26 68.48 61.93 (±0.16) 44.57 44.88 39.82 65.84 51.55 49.33 (±0.37)
LTE 94.96 92.43 85.23 94.19 87.01 90.76 (±1.19) 92.02 80.25 90.30 94.66 90.39 89.52 (±1.47) 47.56 50.73 44.64 97.05 92.90 66.57 (±0.78)
RECIPE 88.27 85.37 84.80 97.78 88.49 88.94 (±0.76) 81.39 70.31 79.19 94.40 95.59 84.18 (±0.74) 47.26 46.30 42.83 100.00 95.38 66.35 (±0.76)
LEMoE 91.59 85.92 86.82 83.64 29.71 75.54 (±0.86) 91.40 90.71 89.12 58.20 49.49 75.79 (±0.48) 87.45 84.54 71.74 85.98 71.23 80.19 (±0.60)
LiveEdit 95.65 93.46 90.98 100.00 100.00 96.02 (±0.93) 98.57 97.92 94.63 100.00 100.00 98.22 (±0.48) 80.65 79.38 75.66 100.00 100.00 87.14 (±0.96)
100 FT-L 52.47 43.68 39.43 91.13 47.72 54.88 (±0.83) 57.26 56.23 56.73 86.19 68.87 65.06 (±0.15) 50.81 48.70 42.79 90.12 52.09 56.90 (±0.89)
FT-M 48.04 40.71 29.86 100.00 33.25 50.37 (±0.43) 56.49 58.16 53.13 100.00 43.52 62.26 (±0.91) 56.23 53.18 43.37 100.00 31.59 56.87 (±0.44)
MEND 17.69 16.39 18.31 91.52 67.94 42.37 (±0.45) 39.91 40.96 40.37 92.32 84.32 59.58 (±0.19) 7.97 8.19 8.04 20.15 23.23 13.52 (±0.13)
SERAC 86.98 81.27 80.97 71.04 15.55 67.17 (±0.60) 66.73 53.69 64.37 59.18 17.86 52.37 (±0.70) 43.55 42.16 39.09 57.77 15.84 39.68 (±0.50)
TP 44.26 38.24 33.28 43.81 38.54 39.63 (±0.10) 46.61 48.12 47.22 64.25 43.21 49.88 (±0.49) 37.51 37.62 33.80 10.26 20.84 28.00 (±0.47)
LTE 92.78 90.36 83.88 94.33 81.79 88.63 (±0.48) 78.41 60.70 78.55 94.34 90.13 80.42 (±0.72) 43.03 42.52 39.63 96.20 91.95 62.67 (±0.81)
RECIPE 88.12 82.49 83.10 98.97 86.53 87.84 (±1.06) 68.79 54.96 67.20 94.61 97.12 76.54 (±0.77) 43.40 42.19 39.02 98.40 95.55 63.71 (±0.86)
LEMoE 29.61 21.70 27.05 79.48 32.61 38.09 (±0.11) 44.13 44.78 42.55 58.13 51.76 48.27 (±0.71) 57.13 51.73 46.73 91.72 58.06 61.08 (±0.34)
LiveEdit 95.25 92.94 85.48 100.00 99.76 94.69 (±0.50) 98.20 97.67 93.96 100.00 100.00 97.97 (±0.34) 79.18 77.44 74.12 100.00 100.00 86.15 (±0.33)
1000 FT-L 45.10 34.62 35.42 48.42 41.24 40.96 (±0.29) 55.39 54.34 53.87 50.80 54.00 53.68 (±0.80) 53.61 48.81 45.92 52.45 59.09 51.98 (±0.70)
FT-M 40.40 31.46 27.85 100.00 27.44 45.43 (±0.68) 47.03 49.68 46.99 100.00 41.41 57.02 (±0.13) 48.24 45.55 43.03 100.00 23.76 52.12 (±0.16)
MEND 15.84 14.35 17.73 91.74 70.17 41.97 (±0.12) 37.22 38.03 37.19 91.49 84.10 57.61 (±0.58) 6.54 6.51 6.50 13.52 20.38 10.69 (±0.14)
SERAC 83.35 70.80 80.32 67.66 13.13 63.05 (±0.87) 53.58 45.78 52.42 56.81 16.90 45.10 (±0.38) 43.12 41.69 38.72 48.08 14.88 37.30 (±0.27)
TP 20.63 15.09 18.41 8.65 8.25 14.21 (±0.18) 24.36 24.21 24.25 16.37 19.96 21.83 (±0.14) 26.03 26.31 24.85 4.11 11.77 18.62 (±0.12)
LTE 89.32 82.82 81.51 94.86 69.83 83.67 (±1.05) 61.67 51.05 61.60 94.78 90.94 72.01 (±0.66) 44.52 43.56 41.29 96.45 90.86 63.33 (±0.34)
RECIPE 84.99 74.20 82.04 96.82 87.73 85.16 (±1.32) 54.64 46.54 54.10 94.60 96.93 69.37 (±1.04) 43.02 41.63 38.59 99.68 92.96 63.18 (±0.16)
LEMoE 19.73 17.34 18.22 72.01 31.06 31.67 (±0.14) 34.74 33.43 32.05 55.55 50.04 41.16 (±0.58) 43.46 43.34 37.69 93.27 67.52 57.06 (±0.38)
LiveEdit 94.42 91.98 84.65 100.00 97.38 93.69 (±0.67) 97.00 91.92 87.53 100.00 100.00 95.29 (±1.48) 72.86 70.34 67.92 100.00 100.00 82.22 (±0.62)
Table 5: Lifelong editing performance on BLIP2-OPT (2.7B) and LLaVA-V1.5 (7B) across the E-VQA and VLKEB datasets.
# Edit Editors E-VQA VLKEB E-IC
Rel. T-Gen. M-Gen. T-Loc. M-Loc. Average Rel. T-Gen. M-Gen. T-Loc. M-Loc. Average Rel. T-Gen. M-Gen. T-Loc. M-Loc. Average
1 FT-L 93.85 86.25 88.58 99.47 84.50 90.53 (±0.20) 82.17 81.64 78.45 98.45 75.06 83.16 (±0.99) 60.40 59.47 52.79 99.95 92.46 73.01 (±1.00)
FT-M 91.42 85.82 49.09 100.00 86.72 82.61 (±0.51) 93.51 91.28 68.30 100.00 81.78 86.97 (±1.36) 71.60 67.34 49.55 100.00 91.69 76.04 (±0.71)
MEND 93.47 91.32 91.22 81.51 74.97 86.50 (±1.17) 90.28 89.18 89.55 93.63 86.69 89.87 (±0.78) 90.28 83.91 79.05 95.92 93.97 88.63 (±0.89)
SERAC 87.66 69.24 85.96 82.11 24.11 69.81 (±0.84) 89.28 89.06 87.35 64.88 12.72 68.66 (±0.76) 85.09 78.96 81.70 83.36 24.50 70.72 (±0.54)
TP 52.45 48.39 51.56 93.93 83.61 65.99 (±0.91) 49.17 51.88 48.92 90.63 79.79 64.08 (±0.98) 52.19 51.49 49.81 83.33 72.04 61.77 (±0.78)
LTE 93.54 77.59 84.07 94.03 88.14 87.48 (±0.45) 92.92 85.16 86.25 88.29 81.96 86.91 (±0.71) 91.30 84.62 83.55 94.96 91.88 89.26 (±1.01)
RECIPE 89.23 70.39 87.94 95.49 90.90 86.79 (±0.42) 93.20 92.40 90.71 90.29 84.35 90.19 (±0.29) 81.25 74.44 76.96 97.18 92.17 84.40 (±0.91)
LEMoE 92.59 90.49 89.16 97.01 84.33 90.71 (±0.94) 92.26 89.28 88.27 93.87 67.49 86.23 (±0.89) 90.32 86.30 85.75 96.66 69.17 85.64 (±1.42)
LiveEdit 94.90 92.63 92.06 100.00 100.00 95.92 (±0.36) 96.41 95.56 92.85 100.00 100.00 96.96 (±1.04) 86.84 83.82 76.95 100.00 100.00 89.52 (±1.20)
10 FT-L 84.14 76.41 71.19 97.84 53.98 76.71 (±0.73) 84.20 81.22 82.12 95.47 68.37 82.28 (±0.73) 62.28 59.67 52.88 98.07 55.84 65.75 (±0.83)
FT-M 88.05 81.83 53.26 100.00 53.01 75.23 (±0.58) 91.31 88.82 63.64 100.00 63.40 81.43 (±1.15) 79.04 70.57 53.82 100.00 53.19 71.33 (±0.31)
MEND 37.34 31.66 35.52 44.31 37.24 37.21 (±0.18) 57.01 55.39 55.34 72.03 67.66 61.49 (±0.47) 69.98 69.35 59.25 93.39 91.52 76.69 (±1.20)
SERAC 86.61 67.86 83.14 67.69 14.47 63.95 (±0.42) 79.90 71.80 78.44 52.64 13.08 59.17 (±0.41) 50.72 49.60 48.36 62.87 16.01 45.51 (±0.47)
TP 35.71 30.29 36.38 74.53 55.67 46.52 (±0.34) 44.70 47.40 44.74 69.61 70.01 55.29 (±0.13) 47.02 47.29 46.06 50.10 42.77 46.65 (±0.64)
LTE 88.55 73.87 83.91 89.44 88.02 84.76 (±0.25) 89.10 81.76 85.96 85.53 81.06 84.68 (±1.26) 54.77 55.69 52.24 90.84 93.06 69.32 (±1.04)
RECIPE 88.44 69.59 85.61 90.68 89.78 84.82 (±1.27) 82.58 74.01 81.18 88.33 85.84 82.39 (±0.20) 50.38 48.96 48.11 92.27 92.69 66.48 (±0.78)
LEMoE 92.55 86.51 83.31 79.76 42.61 76.94 (±0.65) 89.02 86.94 85.18 67.60 50.75 75.90 (±1.13) 88.13 84.15 75.86 92.84 47.92 77.78 (±0.76)
LiveEdit 94.06 91.84 90.60 100.00 99.92 95.28 (±0.97) 95.75 95.13 91.99 100.00 100.00 96.57 (±1.50) 85.79 83.52 76.09 100.00 100.00 89.08 (±1.33)
100 FT-L 64.81 58.11 54.23 93.62 44.98 63.15 (±1.06) 69.98 69.05 68.09 88.74 63.07 71.79 (±0.16) 62.35 57.49 54.13 96.15 46.86 63.39 (±0.17)
FT-M 66.06 51.17 47.24 100.00 36.65 60.22 (±0.28) 60.03 60.18 57.00 100.00 51.24 65.69 (±0.54) 58.93 54.35 47.84 100.00 36.80 59.59 (±0.14)
MEND 23.32 20.75 21.42 47.27 42.60 31.07 (±0.18) 38.85 40.02 37.99 72.82 68.89 51.71 (±0.39) 51.92 53.96 48.24 86.02 83.88 64.80 (±0.79)
SERAC 87.39 66.53 81.70 65.13 14.04 62.96 (±0.64) 69.38 58.31 67.57 51.71 13.72 52.14 (±0.78) 47.41 46.17 45.22 49.37 13.52 40.34 (±0.55)
TP 19.46 20.32 22.03 41.85 32.91 27.32 (±0.39) 39.60 42.52 39.14 36.15 43.68 40.21 (±0.22) 39.60 39.64 38.48 17.29 26.04 32.21 (±0.40)
LTE 85.05 71.80 79.61 87.06 86.10 81.92 (±0.50) 79.45 63.55 79.46 84.18 79.42 77.21 (±1.26) 50.47 49.57 48.17 88.12 92.87 65.84 (±0.70)
RECIPE 88.26 67.74 81.48 91.77 88.23 83.49 (±1.18) 71.54 60.04 70.35 85.57 86.85 74.87 (±1.06) 47.27 46.06 45.12 89.52 94.25 64.44 (±0.68)
LEMoE 27.84 26.38 26.73 68.45 33.73 36.63 (±0.21) 44.75 43.33 41.96 73.90 60.38 52.86 (±0.62) 52.85 50.20 48.11 92.08 49.63 58.57 (±0.53)
LiveEdit 93.23 91.27 88.34 100.00 99.61 94.49 (±0.26) 94.06 94.62 89.42 100.00 100.00 95.62 (±1.26) 84.38 83.53 71.85 100.00 100.00 87.95 (±0.86)
1000 FT-L 58.97 46.34 50.77 72.04 41.92 54.01 (±0.84) 61.85 60.82 60.41 76.17 60.34 63.92 (±0.45) 58.88 53.98 52.98 93.31 46.49 61.13 (±0.74)
FT-M 51.18 42.79 40.08 100.00 37.28 54.27 (±0.76) 51.26 54.35 51.13 100.00 50.72 61.49 (±1.03) 52.65 47.90 47.25 100.00 33.69 56.30 (±0.58)
MEND 31.84 26.60 33.98 43.08 44.90 36.08 (±0.15) 42.35 45.07 42.68 66.98 62.32 51.88 (±0.69) 49.30 51.46 46.70 81.03 82.53 62.20 (±0.30)
SERAC 84.50 60.36 81.83 63.26 13.40 60.67 (±0.22) 57.02 51.66 56.80 48.48 13.24 45.44 (±0.39) 47.03 45.79 44.82 42.12 15.31 39.01 (±0.56)
TP 9.24 8.96 10.25 20.54 17.03 13.20 (±0.19) 23.84 25.00 23.54 18.79 25.58 23.35 (±0.06) 25.07 24.50 24.09 14.77 18.05 21.30 (±0.30)
LTE 81.93 66.81 75.96 88.24 78.95 78.38 (±0.56) 64.28 56.91 64.07 83.96 78.59 69.56 (±0.51) 51.31 49.25 49.19 88.19 90.53 65.69 (±0.91)
RECIPE 85.33 61.79 76.35 90.33 89.08 80.58 (±1.12) 58.18 52.31 58.14 85.49 83.84 67.59 (±0.29) 47.02 45.76 44.85 89.80 92.26 63.94 (±0.79)
LEMoE 25.07 23.87 26.68 89.11 50.50 43.04 (±0.29) 39.48 38.24 36.52 82.35 58.01 50.92 (±0.12) 50.61 47.35 47.19 95.11 47.62 57.58 (±0.91)
LiveEdit 92.75 89.81 85.12 100.00 96.82 92.90 (±0.21) 93.41 89.87 85.21 100.00 100.00 93.70 (±0.31) 81.43 80.34 62.25 100.00 99.73 84.75 (±1.42)
Table 6: Lifelong editing performance on MiniGPT-4 (7B) across the E-VQA and VLKEB datasets.